Two cars leave town going opposite directions. One car is traveling 55 mph, and the other is traveling 65 mph How long will it take before they are 180 miles apart?
Answer by checkley75(3666)

55T+65T=180
120T=180
T=180/120
T=1.5 HOURS THEY WILL BE 180 MILES APART.
PROOF
55*1.5=65*1.5=180
82.5+97.5=180
180=180